OpenStudy (anonymous):

help! how do figure this out A plane can fly 300 miles in the same time as it takes a cr to go 120 miles. If the car travels90 mph slower than the plane, find the speed of the plane

OpenStudy (sandra):

ah ok. so you have some pieces of information that you need to turn into an equation you can solve

OpenStudy (sandra):

I'll give it a shot

OpenStudy (anonymous):

thanks

OpenStudy (sandra):

so we have two basic variables, the speed of the plane, and the speed of the car

OpenStudy (sandra):

since the plane goes 320 miles in the same time the car goes 120 miles, the plane's speed (p) in relation to the car's speed is 320/120

OpenStudy (sandra):

that is: 1. p = 320/120 c

OpenStudy (sandra):

the second piece of information is that the car travels 90 mph slower than the plane

OpenStudy (sandra):

2. c = p - 90

OpenStudy (sandra):

now that you have two equations, and two variables, you can solve via subsitution

OpenStudy (sandra):

so for our case, let's take the value of (c) from the second equation, and substitute it for the (c) in the first equation

OpenStudy (sandra):

1. p = 320/120 c , if c = p - 90, is the same as: 2. p = (320 x (p - 90))/120 3. p = 8(p - 90)/3 (reducing the fraction part) 4. p = 8p - 720/3 5. 3p = 8p - 720 6. 5p = 720 7. p = 144

OpenStudy (sandra):

so the plane travels 144mph